# Rhaly operators acting on Hardy and Bergman spaces

### v1

@@ 4. Rhaly operators acting on Bergman spaces
$\left|\delta_{k+1,N}-\delta_{k,N}\right|\,=\,\frac{\left|\gamma_{k+1,N}-\gamma_{k,N}\right|}{\left|\gamma_{k+1,N}\gamma_{k,N}\right|}\asymp\frac{\left|\alpha_{k+1,N}-\alpha_{k,N}\right|}{\left|\alpha_{k+1,N}\alpha_{k,N}\right|}N^{-\frac{1+\alpha}{p}}.$
This and (3.9) give
(4.4) $\left|\frac{\delta_{k+1,N}-\delta_{k,N}}{\frac{k+1}{N}-\frac{k}{N}}\right|\,\lesssim N^{1-\frac{2+\alpha}{p}},\quad N\leq k\leq 2N.$
For each $N$, let $\Psi_{N}:\mathbb{R}\rightarrow\mathbb{R}$ be defined as follows:
  • $\Psi_{N}(x)=0$, if $x\leq 0$.
  • $\Psi_{N}(x)=0$, if $x\geq 4$.
  • Between 0 and $4$, the graph of $\Psi_{N}$ is the polygonal whose vertices are the points
    $(0,0),\,\left(\frac{N}{N},\delta_{N,N}\right),\,\left(\frac{N+1}{N},\delta_{N+1,N}\right),\,\left(\frac{N+2}{N},\delta_{N+2,N}\right),\,\dots,\left(\frac{2N}{N},\delta_{2N,N}\right),\,(4,0).$
Using (4.4), we see that $\Psi_{N}$ is a Lipschitz function and that
(4.5) $\mathcal{L}(\Psi_{N})\,\lesssim N^{1-\frac{2+\alpha}{p}}.$
Using the results obtained in section 2.3 with $\Psi=\Psi_{N}$ and $n=N$ we see that
$\left|\sum_{k=0}^{4N}\Psi_{N}\left(\frac{k}{N}\right)e^{ik\theta}\right|\lesssim N^{1-\frac{2+\alpha}{p}}\frac{1}{|1-e^{i\theta}|^{2}},\quad 0<\theta<2\pi.$
Hence, if we set
$H_{N}(z)\,=\,\sum_{k=0}^{4N}\Psi_{N}\left(\frac{k}{N}\right)z^{k},\quad z\in\mathbb{D},$
we have that
(4.6) $\|H_{N}\|_{H^{p}}\,\lesssim N^{1-\frac{2+\alpha}{p}}.$
Arguing as in the proof of Theorem 1, using the Riesz projection theorem, we obtain
(4.7) $\|\Delta_{N}\left(F_{(\eta)}^{\prime}\right)\|_{A^{p}_{\alpha}}\lesssim\|H_{N}\ast\mathcal{R}_{(\eta)}(g_{N})\|_{A^{p}_{\alpha}}.$
Since $\mathcal{R}_{(\eta)}\in\mathcal{B}(A^{p}_{\alpha})$, we have $\|\mathcal{R}_{(\eta)}(g_{n})\|_{A^{p}_{\alpha}}\lesssim 1$ and then, using Lemma 2 and (4.6), it follows that
$\|\Delta_{N}\left(F_{(\eta)}^{\prime}\right)\|_{A^{p}_{\alpha}}\,\lesssim N^{1-\frac{2+\alpha}{p}}.$
Using Lemma 3. 1 of [29] it follows easily that $\|\Delta_{N}\left(F_{(\eta)}^{\prime}\right)\|_{H^{p}}\asymp\,N^{\frac{1+\alpha}{p}}\|\Delta_{N}\left(F_{(\eta)}^{\prime}\right)\|_{A^{p}_{\alpha}}$. Then it follows that $\|\Delta_{N}\left(F_{(\eta)}^{\prime}\right)\|_{H^{p}}\lesssim N^{1-\frac{1}{p}}$, which is equivalent to saying that $F_{(\eta)}\in\Lambda^{p}_{1/p}$.
Suppose now that $\mathcal{R}_{(\eta)}\in\mathcal{K}(A^{p}_{\alpha})$. Since $g_{N}\,\rightarrow\,0$, as $N\to\infty$, uniformly in compact subsets, it follows that $\mathcal{R}_{(\eta)}(g_{N})\,\rightarrow\,0$, as $N\to\infty$, in $A^{p}_{\alpha}$. This leads to $\|\Delta_{N}\left(F_{(\eta)}^{\prime}\right)\|_{A^{p}_{\alpha}}\,=\operatorname{o}\left(N^{1-\frac{2+\alpha}{p}}\right)$ and to $\|\Delta_{N}\left(F_{(\eta)}^{\prime}\right)\|_{H^{p}}=\,\operatorname{o}\left(N^{1-\frac{1}{p}}\right)$. Hence, $F_{(\eta)}\in\lambda^{p}_{1/p}$. $\Box$
5. Further results
@@
$\Lambda^{1}_{1}\,=\,\{f\in\rm{Hol}(\mathbb{D}):f^{\prime}\in H^{1}\}.$
@@
It is well known that there are function $h$ in $H^{1}$ for which the sequence of partial sums $S_{N}h$ does not converge. However, the following is true:
If $h\in H^{1}$ then there exists a subsequence $\{S_{N_{k}}h\}$ of the sequence of partial sums of $h$ such that $S_{N_{k}}h\to f$, as $k\to\infty$, in the norm of $H^{1}$.
@@
$\frac{1}{N}\sum_{n=1}^{N}\frac{|\eta_{n}|}{n+1}\left(\sum_{k=1}^{n}ka_{N}^{k}\right)\leq\frac{1}{N}\sum_{n=1}^{\infty}\frac{|\eta_{n}|}{n+1}\left(\sum_{k=1}^{n}ka_{N}^{k}\right)\leq\pi\|\mathcal{R}_{(\eta)}(f_{a})\|_{H^{1}}\,\lesssim 1.$
@@
$\Delta_{N}(F_{(\eta)})\,=\,N\mathcal{R}_{(\eta)}(f_{N})\ast\varphi_{N}.$
Then it follows that
(5.3) $\|\Delta_{N}(F_{(\eta)})\|_{H^{1}}\,\lesssim N\|\mathcal{R}_{(\eta)}(f_{N})\|_{H^{1}}\|\varphi_{N}\|_{H^{1}}\lesssim N\|\psi_{N}\|_{H^{1}}.$
Notice that the sequence of Taylor coefficients of $\psi_{N}$ is decreasing. Then using [33, Theorem 1. 1], we obtain
$\|\psi_{N}\|_{H^{1}}\,\asymp\,\sum_{n=0}^{N-1}\frac{1}{(n+1)\sum_{k=1}^{n+N}ka_{N}^{k}}\,\leq\,\sum_{n=0}^{N-1}\frac{1}{(n+1)\sum_{k=1}^{N}ka_{N}^{k}}\,\asymp\,\frac{\log N}{N^{2}}.$
This and (5.3) shows that $\|\Delta_{N}(F_{(\eta)})\|_{H^{1}}\,\lesssim\,\frac{\log N}{N}$. Then, using Lemma 3. 4 of [10], we obtain
@@
$\Upsilon(z)\,=\,\sum_{k=0}^{\infty}\left(\Delta_{2_{k}}h\right)_{t_{k}}(z),\quad z\in\mathbb{D}.$
Clearly, $\Upsilon\in\rm{Hol}(\mathbb{D})$ and $\Delta_{2^{k}}\Upsilon=\left(\Delta_{2^{k}}h\right)_{t_{k}}$, for all $k$. Let $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ be defined by $\eta_{n}=\frac{\Upsilon^{(n}(0)}{n!}$. Then $\Upsilon=F_{(\eta)}$. We have that $|\eta_{n}|=\frac{1}{n}$ for all $n\geq 1$. Also
@@
If $1<p<2$, then (5.5) implies that $F_{(\eta)}\not\in\Lambda^{p}_{1/p}$ and then, using Theorem 1, we deduce that $\mathcal{R}_{(\eta)}\not\in\mathcal{B}(H^{p})$.
@@
Part (iii) of Theorem 4 follows immediately from the following result (see [20] for $p\geq 2$, and [30, Lemma 1] for $1<p<\infty$).

### v2

@@ 5. Sufficient conditions
$M_{1}(r^{2},\phi)\,\lesssim M_{\gamma p}(r^{2},f)\left(\int_{0}^{2\pi}\frac{d\theta}{|1-re^{i\theta}|^{2\sigma}}\right)^{1/\sigma}\asymp\frac{1}{(1-r)^{1+\frac{1}{\gamma p}}}M_{\gamma p}(r^{2},f)$
and then it follows that
$\int_{0}^{1}(1-r)^{\alpha+1}M_{1}^{p}(r^{2},\phi)\,dr\,\lesssim\int_{0}^{1}(1-r)^{\alpha+1-p-\frac{1}{\gamma}}M_{\gamma p}^{p}(r^{2},f)\,dr.$
Since $\alpha+1-p-\frac{1}{\gamma}>-1$, using Theorem 5. 6 of [12], we obtain
$\int_{0}^{1}(1-r)^{\alpha+1}M_{1}^{p}(r^{2},\phi)\,dr\,\lesssim\int_{0}^{1}(1-r)^{\alpha+1-\frac{1}{\gamma}}M_{\gamma p}^{p}(r^{2},f^{\prime})\,dr.$
Now, a slight modification of the proof of Theorem 5. 9 of [12] gives that
$M_{\gamma p}(r^{2},f^{\prime})\lesssim\frac{1}{(1-r)^{\frac{1}{p}-\frac{1}{\gamma p}}}M_{p}(r,f^{\prime}).$
Then we obtain
$\int_{0}^{1}(1-r)^{\alpha+1}M_{1}^{p}(r^{2},\phi)\,dr\,\lesssim\int_{0}^{1}(1-r)^{\alpha}M_{p}^{p}(r,f^{\prime})\,dr\,\lesssim\|f\|_{\mathcal{D}^{p}_{\alpha}}^{p}.$
This, (5.17), and (5.16) give
(5.18) $\int_{0}^{1}(1-r)^{\alpha+p}M_{p}^{p}(r^{4},\mathcal{R}(f)^{\prime\prime})\,dr\,\lesssim\beta_{p,\frac{1}{p}}^{p}(F_{(\eta)})\|f\|_{\mathcal{D}^{p}_{\alpha}}^{p}.$
Hence (5.15) holds as desired.
Suppose now that $1<p<\infty$ and $F_{(\eta)}\in\lambda^{p}_{1/p}$. For $N=2,3,4,\dots$ let $\mathcal{R}_{N}$ be defined as in (5.11). Using (5.18) and arguing as in the proof of Proposition 3, we deduce, for $f\in\mathcal{D}^{p}_{\alpha}$,
$\left\|\left(\mathcal{R}_{(\eta)}\,-\,\mathcal{R}_{N}\right)(f)\right\|_{\mathcal{D}^{p}_{\alpha}}^{p}\lesssim\beta^{p}_{p,\frac{1}{p}}\left(F_{(\eta)}\,-\,S_{N}F_{(\eta)}\right)\|f\|_{\mathcal{D}^{p}_{\alpha}}^{p}.$
Using Proposition 1 we deduce that $\{\mathcal{R}_{(\eta)}-\mathcal{R}_{N}\}\,\rightarrow\,0$, as $N\to\infty$, in the norm of $\mathcal{B}(\mathcal{D}^{p}_{\alpha})$. Since the operators $\mathcal{R}_{N}:\mathcal{D}^{p}_{\alpha}\rightarrow\mathcal{D}^{p}_{\alpha}$ are finite rank operators, it follows that $\mathcal{R}_{(\eta)}\in\mathcal{K}(\mathcal{D}^{p}_{\alpha})$. Thus, (ii) is proved.
$\Box$
6. Rhaly operators acting on $H^{1}$ and some further results
@@
$\Lambda^{1}_{1}\,=\,\{f\in\rm{Hol}(\mathbb{D}):f^{\prime}\in H^{1}\}\,=\,S^{1}.$
@@
It is well known that there are functions $h$ in $H^{1}$ for which the sequence of partial sums $S_{N}h$ does not converge. However, the following is true:
If $h\in H^{1}$ then there exists a subsequence $\{S_{N_{k}}h\}$ of the sequence of partial sums of $h$ such that $S_{N_{k}}h\to h$, as $k\to\infty$, in the norm of $H^{1}$.
@@
$\frac{1}{N}\sum_{n=1}^{N}\frac{|\eta_{n}|}{n+1}\left(\sum_{k=1}^{n}ka_{N}^{k}\right)\leq\pi\|\mathcal{R}_{(\eta)}(f_{N})\|_{H^{1}}\,\lesssim 1.$
@@
$\Delta_{N}(F_{(\eta)})\,=\,N\mathcal{R}_{(\eta)}(f_{N})\ast\varphi_{N}$
where $\varphi_{N}$ is defined in (4.1). Using this, (6.1), and Proposition 2, we obtain that $\|\Delta_{N}(F_{(\eta)})\|_{H^{1}}\,\lesssim\,\frac{\log N}{N}$. Then, using Lemma 3. 4 of [11], we obtain
@@
$\Upsilon(z)\,=\,\sum_{k=0}^{\infty}\left(\Delta_{2^{k}}h\right)_{t_{k}}(z),\quad z\in\mathbb{D}.$
Clearly, $\Upsilon\in\rm{Hol}(\mathbb{D})$ and $\Delta_{2^{k}}\Upsilon=\left(\Delta_{2^{k}}h\right)_{t_{k}}$, for all $k$. Let $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ be defined by $\eta_{n}=\frac{\Upsilon^{(n)}(0)}{n!}$. Then $\Upsilon=F_{(\eta)}$. We have that $|\eta_{n}|=\frac{1}{n}$ for all $n\geq 1$. Also
@@
If $1<p<2$, then (6.4) implies that $F_{(\eta)}$ does not satisfy (1.2) and then, using Theorem 1, we deduce that $\mathcal{R}_{(\eta)}\not\in\mathcal{B}(H^{p})$.
@@
Part (iii) of Theorem 5 follows immediately from the following result (see [22] for $p\geq 2$, and [32, Lemma 1] for $1<p<\infty$).
@@
7. Rhaly operators acting on derivative-Hardy spaces
The proof of Theorem 6 will be based on the Littlewood-Paley characterization of Hardy spaces. Namely, we shall use the following result which can be found in [46, Vol. II, Chapter XIV, Theorem 4. 14].
Theorem C.
Suppose that $1<q<\infty$ and let $\{\lambda_{j}\}_{j=0}^{\infty}$ be a sequence of complex numbers. Let $M$ be a positive number and suppose that
$|\lambda_{j}|\leq M,\,\,\,\,\sum_{j=2^{k}}^{2^{k+1}-1}|\lambda_{j+1}-\lambda_{j}|\,\leq M,\quad j,k=0,1,2,\dots.$
There exists a positive constant $A_{q}$ such that if $f\in H^{q}$, $f(z)=\sum_{n=0}^{\infty}a_{n}z^{n}$ ($z\in\mathbb{D}$), and
$h(z)=\sum_{n=0}^{\infty}\lambda_{n}a_{n}z^{n},\quad z\in\mathbb{D},$
then $h\in H^{q}$ and
$\|h\|_{H^{q}}\,\leq\,MA_{q}\|f\|_{H^{q}}.$
Proof ofTheorem 6.
The implication (i)  $\Rightarrow$   (iii) is clear. Indeed,
$\mathcal{R}_{(\eta)}\in\mathcal{B}(S^{p},S^{q})\,\,\Rightarrow\,\,\mathcal{R}_{(\eta)}(1)\in S^{q}\,\,\Leftrightarrow\,\,F_{(\eta)}\in S^{q}.$
Let us prove next that (iii)  $\Rightarrow$   (i). Assume that $F_{(\eta)}\in S^{q}$, that is, $F_{(\eta)}^{\prime}\in H^{q}$. Take $f\in S^{p}$, $f(z)=\sum_{n=0}^{\infty}a_{n}z^{n}$ ($z\in\mathbb{D}$). Then $f^{\prime}\in H^{1}$ and then, using Hardy’s inequality, we see that
$\sum_{n=0}^{\infty}|a_{n+1}|\leq\pi\|f^{\prime}\|_{H^{1}}\leq\pi\|f^{\prime}\|_{H^{p}}.$
This implies that
(7.1) $\sum_{n=0}^{\infty}|a_{n}|\leq\pi\|f\|_{S^{p}}.$
We have
$\mathcal{R}_{(\eta)}(f)(z)=\sum_{j=0}^{\infty}\eta_{j}\left(\sum_{k=0}^{j}a_{k}\right)z^{j},\quad z\in\mathbb{D}.$
Hence,
$\mathcal{R}_{(\eta)}(f)^{\prime}(z)=\sum_{j=0}^{\infty}(j+1)\eta_{j+1}\left(\sum_{k=0}^{j+1}a_{k}\right)z^{j},\quad z\in\mathbb{D}.$
Set
$\lambda_{j}\,=\,\sum_{k=0}^{j+1}a_{k},\quad j=0,1,2,\dots.$
Using (7.1), se see that
$|\lambda_{j}|\leq\sum_{k=0}^{j+1}|a_{k}|\leq\sum_{k=0}^{\infty}|a_{k}|\leq\pi\|f\|_{S^{p}},$
$\sum_{j=0}^{\infty}|\lambda_{j+1}-\lambda_{j}|=\sum_{j=0}^{\infty}|a_{j+2}|\leq\pi\|f\|_{S^{p}}.$
Then Theorem C yields that
$\|\mathcal{R}_{(\eta)}(f)^{\prime}\|_{H^{q}}\leq A_{q}\pi\|f\|_{S^{p}}\|F_{(\eta)}^{\prime}\|_{H^{q}}.$
Thus (i) follows.
The implication (ii)  $\Rightarrow$  (i) is obvious. Let us prove that (iii)  $\Rightarrow$  (ii). So, assume (iii), that is, $F_{(\eta)}\in S^{q}$. Then we already know that $\mathcal{R}_{(\eta)}\in\mathcal{B}(S^{p},S^{q})$.
Since $1<q<\infty$, we have that
(7.2) $S_{N}(F_{(\eta)}^{\prime})\to F_{(\eta)}^{\prime},\quad\text{as $N\to\infty$,\,\,in $H^{q}$}.$
For $N=2,3,4,\dots$, let $R_{N}$ be defined as above, that is, if $f\in S^{p}$, $f(z)=\sum_{n=0}^{\infty}a_{n}z^{n}$ ($z\in\mathbb{D}$), then
$R_{N}(f)(z)\,=\,\sum_{j=0}^{N}\eta_{j}\left(\sum_{k=0}^{j}a_{k}\right)z^{j},\quad z\in\mathbb{D}.$
Then
$\mathcal{R}_{(\eta)}(f)^{\prime}(z)-R_{N}(f)^{\prime}(z)\,=\,\sum_{j=N}^{\infty}(j+1)\eta_{j+1}\left(\sum_{k=0}^{j+1}a_{k}\right)z^{j}.$
Set
$\lambda_{j}\,=\,\left\{\begin{array}[]{ll}0,&\hbox{if $j=0,1,2,\dots N$}.\\ \\ \sum_{k=0}^{j+1}a_{k},&\hbox{if $j=N,N+1,\dots$}.\end{array}\right$.
Just as above we have that
$|\lambda_{j}|\leq\pi\|f\|_{S^{p}},\,\,\text{for all $j$,\,\, and}\,\,\,\,\sum_{j=0}^{\infty}|\lambda_{j+1}-\lambda_{j}|\leq\pi\|f\|_{S^{p}}.$
Then Theorem C implies that
$\|\mathcal{R}_{(\eta)}(f)^{\prime}(z)-R_{N}(f)^{\prime}(z)\|_{H^{q}}\leq A_{q}\pi\|F_{(\eta)}^{\prime}-S_{N-1}(F_{(\eta)}^{\prime})\|_{H^{q}}\|f\|_{S^{p}}.$
This and (7.2) give that $R_{N}\to\mathcal{R}_{(\eta)}$ in $\mathcal{B}(S^{p},S^{q})$. Since the operators $R_{N}$ are finite rank operators from $S^{p}$ into $S^{q}$, it follows that $\mathcal{R}_{(\eta)}\in\mathcal{K}(S^{p},S^{q})$. $\Box$